OpenStudy (anonymous):

Write an equation in point-slope form for the line through the given point with the given slope. (8, 10); m = 5/2 a. y – 10 = (5/2)(x – 8) b. y – 8 = (5/2)(x – 10) c. y + 10 = (5/2)(x + 8) d. y + 10 = (5/2)(x – 8)

OpenStudy (anonymous):

point slope form is y - yo = m ( x - xo) plug in y-10 = 5/2 ( x-8) That is A